### v1

@@ Abstract
Divisor methods are well known to satisfy house monotonicity, which allows representative seats to be allocated sequentially. We focus on stationary divisor methods defined by a rounding cut point $c\in[0,1]$. For such methods with integer-valued votes, the resulting apportionment sequences are periodic. Restricting attention to two-party allocations, we characterize the set of possible sequences and establish a connection between the lexicographical ordering of these sequences and the parameter $c$. We then show how sequences for all pairs of parties can be systematically extended to the $n$-party setting. Further, we determine the number of distinct sequences in the $n$-party problem for all $c$. Our approach offers a refined perspective on large-party bias: rather than viewing large parties as simply receiving more seats, we show that they instead obtain their seats earlier in the apportionment sequence. Of particular interest is a new relationship we uncover between the sequences generated by the smallest divisors (Adams) and greatest divisors (d’Hondt or Jefferson) methods.
@@
apportionment sequence, divisor methods, d’Hondt/Jefferson method, Adams method
@@ 1 Introduction
Divisor methods of apportionment are widely used in proportional representation systems to allocate seats to political parties in proportion to their share of the popular vote. Less attention, however, has been given to how these methods determine the sequence in which seats are awarded. This sequence is important in practice as it is used by some coalition governments to determine cabinet positions: while the office of prime minister typically goes to the party with a plurality of votes, other cabinet positions are selected by parties based on the sequence. For example, in Northern Ireland, the greatest divisors method (also known as d’Hondt’s or Jefferson’s method) is used not only to determine the number of seats allocated to each party in the Assembly but also the order in which cabinet posts are selected [14]. Sequential apportionment in this context offers an alternative to the intense negotiation and strategic bargaining that often arise during the establishment of a new cabinet. Using an apportionment method to allocate positions sequentially can help mitigate these conflicts and avoid delays.
@@
Our work has application to sequential apportionment in both fields, as we focus our attention on the possible apportionment sequences from stationary divisor methods, a family parametrized by a rounding cut point $c\in[0,1]$ which includes the smallest divisors method (also known as Adams’ method for which $c=0$), Webster’s or Sainte Laguë’s method ($c=0.5$), and d’Hondt’s method (used in Northern Ireland, for which $c=1$). Such sequences arise whenever apportionment methods satisfy the property of house monotonicity, which requires that no party’s allocation decreases as the number of total available seats increases. Monotonic apportionment methods give rise naturally to a sequence by considering the allocation of one seat (the first seat), the allocation of the second seat, and so on. Stationary divisor methods have an easily described algorithm for determining this sequence—versions of which are used in awarding minister positions and in JIT sequencing problems.
@@
In Section 3, we focus our attention on properties of 2-party apportionment sequences. For any vote distribution, we characterize the possible 2-party apportionment sequences for all stationary divisor methods, breaking the interval of possible cut points $[0,1]$ into equivalence classes. The possible apportionment sequences for $n$ parties can be determined by lifting all $\binom{n}{2}$ pairs of 2-party sequences to the $n$-party sequence. Using properties of the common divisors of the parties’ vote totals, we are able to count the number of possible sequences for the $n$-party problem, discussed in Section 4. We offer some concluding remarks in Section 5.
@@ 2 Sequential Apportionment
Proportional representation systems assign seats to parties in roughly the same proportion as the votes they receive. The d’Hondt method—also called the greatest divisors method—assigns seats to each party sequentially by dividing their vote count by successive divisors $1,2,3,\dots$ and allocating the next seat to the party with the greatest ratio. At each step, the divisor for the party that received the previous seat is updated. The process is repeated until all seats have been allocated. We illustrate with an example from the 2022 Northern Ireland Assembly. This election was notable in part because it was the first election in which Sinn Féin was awarded the largest number of seats. Eight parties received seats, based on the number of votes received. Following the election, positions for the Executive Committee were selected sequentially using the d’Hondt method. However, instead of using the popular vote totals, the proportions were based on the number of seats received by each party in the Assembly.
@@
The 2022 election for the Northern Ireland Assembly resulted in the following eight parties receiving seats:
@@
The top two vote-getting parties, Sinn Féin (SF) and the Democratic Unionist Party (DUP), nominated individuals for the First Minister and the deputy First Minister, respectively. The position of Justice Minister was awarded to the Alliance Party (AP) as part of a political compromise requiring cross-community support from both nationalists and unionists. The remaining positions were then allocated to the parties sequentially. For instance, the next position was allocated to SF because
@@
because the divisor for SF had been updated in a similar manner.
@@
$27/1$ >> $25/1$ >> $27/2$ >> $25/2$ >> $9/1$ $\geq$ $27/3$ >> $17/2$
Even though the use of sequential apportionment is supposed to mitigate delays in the negotiation of minister positions, the Democratic Unionist Party still delayed the formation of the Executive Committee, which resulted in legal action to prevent a snap election. Thus, this selection process was not completed until 2024, and the use of sequential apportionment did not in this case decrease the delay in forming the Executive Committee.
The use of the divisors $1,2,3,\ldots,$ distinguishes D’Hondt’s method. More generally, divisor methods can be defined through their unique set of denominators or divisors. In this article, we focus on the class of stationary divisor methods, which are parametrized by a fixed constant $c\in[0,1]$, and given by the sequence $a+c$ where $a$ is the nonnegative integer number of positions already awarded, that is, the current apportionment. We refer to the value $c$ as the cut point of the corresponding divisor method. The process of sequential apportionment for a general cut point $c$ is defined below.
@@
For fixed $0\leq c\leq 1$, given a distribution of votes $\mathbf{p}=(p_{1},p_{2},\ldots,p_{n})$, the apportionment sequence for the stationary divisor method with cut point $c$ is defined inductively as follows. Start by awarding each party $a_{i}=0$ seats. Then if party $p_{i}$ has been awarded $a_{i}$ seats, award the next seat to the party that maximizes the ratio $\frac{p_{i}}{a_{i}+c}$. Continue until all $h$ seats have been allocated.
If $c=1$, then the sequential apportionment process corresponds to the D’Hondt method. If $c=1/2$, the process corresponds to the Sainte-Lagué or Webster method. If $c=0$, the apportionment method is known as Adams’ method. Because $a_{i}=0$ for all parties $i$ before any seats have been allocated, Adams’ method is undefined because $\frac{p_{i}}{a_{i}}$ is undefined, because dividing by zero is undefined. To apply Adams’ method, it is assumed that $\frac{p_{i}}{0}>\frac{p_{j}}{0}$ whenever $p_{i}>p_{j}$ and $\frac{p_{i}}{0}>\frac{p_{j}}{k}$ for any $k>0$. Thus, under Adams’ method, the first $n$ seats are awarded, one to each party, assuming there are sufficient seats. There are other divisor methods, such as the Hill-Huntington method and Dean’s method, based on the geometric and harmonic means, respectively, that also give a seat to every party before a party gets a second seat. (To see how Hill-Huntington’s method fits into the sequential framework, it awards the next seat to the party that maximizes the ratio $p_{i}/\sqrt{a_{i}(a_{i}+1)}$, where $\sqrt{a_{i}(a_{i}+1)}$ is the geometric mean of the current apportionment $a_{i}$ and the next apportionment $a_{i}+1$.)
@@
Suppose two parties receive votes $\mathbf{p}=(32,14)$ with a house size of $h=5$. If $c=1$ then the first two seats are given to party 1 since $32/1>32/2>14/1$. The third and fourth seats are given to party 2 and party 1, respectively, since $14/1>32/3>14/2$. The fifth seat is given to party 2 since $32/4>14/2$.
If $c=0$, then each party receives a seat: party 1 receives the first seat and party 2 receives the second seat. The third and fourth seats are given to $1$ since, again, $32/1>32/2>14/1$. The fifth seat is given to $2$ since $\frac{14}{1}>\frac{32}{3}$.
In the example, notice that after 5 seats have been given out, parties 1 and 2 have received 3 and 2 seats, respectively, under both D’Hondt’s and Adams’ methods, however the sequence with which these seats have been allocated is different. We focus on the apportionment sequences that include apportionments for all house sizes $h$ by considering the infinite sequence.
@@
For fixed $\mathbf{p}$, cut point $c$, and a house size $h$, let $S_{c}(h,\mathbf{p})=(s_{1},s_{2},\dots,s_{h})$ where $s_{j}=i$ if party $i$ receives the $j$th seat for each $j=1,\ldots,h$. Let $S_{c}(\mathbf{p})=\lim_{h\rightarrow\infty}S_{c}(h,\mathbf{p})$.
Thus, in Example 2, $S_{0}(5,\mathbf{p})=(1,\,2,\,1,\,1,\,2)$ and $S_{1}(5,\mathbf{p})=(1,\,1,\,2,\,1,\,2)$. Notice that after the first two seats have been allocated to states 1 and 2 under Adams’ method, the sequence for the next three seats is $1,1,2$, which is the same as the first three seats under d’Hondt’s method. This is because once each state has been awarded a seat, the denominators in each ratio under Adams’ method, $p_{i}/(a_{i}+1),$ are the same as under the d’Hondt method. This relationship holds more generally and leads to the following theorem.
@@
For every $\mathbf{p}$, if $S_{1}(\mathbf{p})=(s_{1},s_{2},\dots)$ is the sequential apportionment under D’Hondt’s method then the sequential apportionment under Adams’ method is $S_{0}(\mathbf{p})=(1,2,\dots,n,s_{1},s_{2},s_{3},\dots)$.
@@
Under Adams’ method, each of the parties $1,2,\dots,n$ receive, in this order, one of the first $n$ seats. After that, each party has one seat, and so the next seat is given to the party maximizing $\frac{p_{i}}{1}$. These are the same ratios whose maximum determines the first seat under D’Hondt’s method; hence, seat $n+1$ is given to party 1, or to party $s_{1}$ (because $p_{1}\geq p_{i}$ for all $i$). The next seat is given to the party corresponding to the maximum of $\{\{\frac{p_{i}}{1}\}_{i\neq s_{i}},\frac{p_{s_{i}}}{2}\}$. Again, these are the same ratios used to determine the second seat under D’Hondt’s method; so it is given to party $s_{2}$, and so on. ∎
@@ 2.1 A Refinement of the Notion of Bias
Many properties of apportionment methods are best described in terms of the number of seats each party receives. Let $F_{c}(h,\mathbf{p})=\mathbf{a}=(a_{1},a_{2},\ldots,a_{n})$ indicate that after $h$ seats, the apportionment method has allocated $a_{i}$ seats to party $p_{i}$, where $h=\sum_{i}a_{i}$. Thus, in Example 2, $F_{0}(5,\mathbf{p})=F_{1}(5,\mathbf{p})=(3,2)$; however $F_{0}(2,\mathbf{p})=(1,1)\neq(2,0)=F_{1}(2,\mathbf{p}).$
When compared to one another, divisor methods are subject to some amount of bias, either favoring large or small parties (based on the number of votes received). The following definition and proposition are due to [2]
@@
If $F_{c}$ and $F^{\prime}_{c^{\prime}}$ are stationary divisor methods corresponding to cut points $c$ and $c^{\prime}$, respectively, with $c>c^{\prime}$, then $F$ favors large parties relative to $F^{\prime}$.
@@
Suppose $c>c^{\prime}$ and that there exists $\mathbf{p}$ and an $h$ such that $S_{c}(h,\mathbf{p})>S_{c^{\prime}}(h,\mathbf{p})$. Let $h^{\prime}+1$ be the first index for which $S_{c}(h,\mathbf{p})$ and $S_{c^{\prime}}(h,\mathbf{p})$ differ. Thus
@@
$cp_{i}+c^{\prime}p_{j}>cp_{j}+c^{\prime}p_{i}\leftrightarrow(c-c^{\prime})(p_{i}-p_{j})>0.$
Since $c>c^{\prime}$, this implies $p_{i}>p_{j}$ and hence $i<j$. This contradiction proves the claim. ∎
Among stationary divisor methods, D’Hondt’s and Adams’ methods are the most and least biased towards large parties, respectively. Hence, $S_{0}(h,\mathbf{p})\geq S_{1}(h,\mathbf{p})$ for every $h$ and $S_{0}(\mathbf{p})\geq S_{1}(\mathbf{p})$ and, by Proposition 3, $S_{0}(h,\mathbf{p})\geq S_{c}(h,\mathbf{p})\geq S_{1}(h,\mathbf{p})$ for all other cut points $c\in(0,1)$. We will use the lexicographical ordering of apportionment sequences to refine the notion of bias, thereby showing how bias is less an absolute statement and more a reflection of when large or small parties receive their seats.
To achieve this refinement of bias, we need to recall additional properties (homogeneity and weak proportionality) of divisor methods. A divisor method satisfies homogeneity because the apportionment only depends on the relative proportion of votes $p_{i}/P$ where $P=p_{1}+\cdots+p_{n}$. Consequently, for a stationary divisor method with cut point $c$, if $\lambda>0$ is a constant then $F_{c}(h,\mathbf{p})=F_{c}(h,\lambda\mathbf{p})$.
Weak proportionality means that if a vector of seats $\mathbf{a}=(a_{1},a_{2},\ldots,a_{n})$ is proportional to $\mathbf{p}=(p_{1},p_{2},\ldots,p_{n})$, then after $h=\sum_{i}a_{i}$ seats have been allocated, party $i$ has received exactly $a_{i}$ seats, or $F_{c}(\sum_{i}a_{i},\mathbf{p})=\mathbf{a}$. We illustrate this with a continuation of Example 2.
@@
Let $\mathbf{p}=(32,14)$. Since $\mathbf{a}=(16,7)$ and $\mathbf{a}=(32,14)$ are both proportional to $\mathbf{p}$, then under D’Hond’t method, $F_{1}(23,\mathbf{p})=(16,7)$ and $F_{1}(46,\mathbf{p})=(32,14).$ Additionally, their corresponding sequences are
@@
The proof that divisor methods satisfy weak proportionality follows most easily from an equivalent (non-sequential) definition of divisor methods. Under this definition, each apportionment method corresponds to a “rounding rule” $f$ that assigns to each positive integer $k$ a value $f(k)\in[k-1,k]$. Numbers in [k1,f(k))[k-1,f(k)) are rounded to k1k-1 and numbers in $[f(k),k]$ are rounded up to $k$. (The function must also satisfy the property that there cannot exist positive integers $a$ and $b$ for which f(a)=a1f(a)=a-1 and $f(b)=b$.) The divisor method associated with rounding rule $f$ is defined as follows.

### v2

@@ Abstract
Divisor methods are well known to satisfy house monotonicity, which allows representative seats to be allocated sequentially. We focus on stationary divisor methods defined by a rounding cutpoint $c\in[0,1]$. For such methods with integer-valued votes, the resulting apportionment sequences are periodic. Restricting attention to two-party allocations, we characterize the set of possible sequences and establish a connection between the lexicographical ordering of these sequences and the parameter $c$. We then show how sequences for all pairs of parties can be systematically extended to the $n$-party setting. Further, we determine the number of distinct sequences in the $n$-party problem for all $c$. Our approach offers a refined perspective on size bias: rather than viewing large parties as simply receiving more seats, we show that they instead obtain their seats earlier in the apportionment sequence. Of particular interest is a new relationship we uncover between the sequences generated by the smallest divisor (Adams) and greatest divisor (D’Hondt or Jefferson) methods.
@@
apportionment sequences, stationary divisor methods, D’Hondt/Jefferson method, Adams method
@@ 1 Introduction
Divisor methods of apportionment are widely used in proportional representation systems to allocate seats to political parties in proportion to their share of the popular vote. Less attention, however, has been given to how these methods determine the sequence in which seats are awarded. This sequence is important in practice as it is used by some coalition governments to determine cabinet positions: while the office of prime minister typically goes to the party with a plurality of votes, other cabinet positions are selected by parties based on the sequence. For example, in Northern Ireland, the greatest divisor method (also known as D’Hondt’s or Jefferson’s method) is used not only to determine the number of seats allocated to each party in the Assembly but also the order in which cabinet posts are selected [18]. Sequential apportionment in this context offers an alternative to the intense negotiation and strategic bargaining that often arise during the establishment of a new cabinet. Using an apportionment method to allocate positions sequentially can help mitigate these conflicts and avoid delays.
@@
Our work has application to sequential apportionment in both fields, as we focus our attention on the possible apportionment sequences from stationary divisor methods, a family parametrized by a rounding cutpoint $c\in[0,1]$ which includes the smallest divisor method (also known as Adams’ method for which $c=0$), Webster’s or Sainte Laguë’s method ($c=0.5$), and D’Hondt’s method (used in Northern Ireland, for which $c=1$). Such sequences arise whenever apportionment methods satisfy the property of house monotonicity, which requires that no party’s allocation decreases as the number of total available seats increases. Monotonic apportionment methods give rise naturally to a sequence by considering the allocation of one seat (the first seat), the allocation of the second seat, and so on. Stationary divisor methods have an easily described algorithm for determining this sequence—versions of which are used in awarding minister positions and in JIT sequencing problems.
@@
In Section 3, we focus our attention on properties of 2-party apportionment sequences. For any vote distribution, we characterize the possible 2-party apportionment sequences for all stationary divisor methods, breaking the interval of possible cutpoints $[0,1]$ into equivalence classes. The possible apportionment sequences for $n$ parties can be determined by lifting all $\binom{n}{2}$ pairs of 2-party sequences to the $n$-party sequence. Using properties of the common divisors of the parties’ vote totals, we are able to count the number of possible sequences for the $n$-party problem. Our analyses for $n$-party sequential apportionment appears in Section 4. We offer some concluding remarks in Section 5. An appendix includes some of the more technical proofs.
@@ 2 Sequential Apportionment
Proportional representation systems assign seats to parties in roughly the same proportion as the votes they receive. The D’Hondt method—also called the greatest divisor method—assigns seats to each party sequentially by dividing their vote count by successive divisors $1,2,3,\dots$ and allocating the next seat to the party with the greatest ratio. At each step, the divisor for the party that received the previous seat is updated. The process is repeated until all seats have been allocated. We illustrate with an example from the 2022 Northern Ireland Assembly. This election was notable in part because it was the first election in which Sinn Féin was awarded the largest number of seats. Eight parties received seats, based on the number of votes received. The First Minister and deputy First Minister were nominated by the largest and second largest parties, respectively [21]. The remaining eight positions of the Executive Committee were selected sequentially using the D’Hondt method. However, instead of using the popular vote totals, the proportions were based on the number of seats received by each party in the Assembly.
@@
The 2022 election for the Northern Ireland Assembly resulted in the following eight parties receiving seats [17]; the parties and their number of seats were:
@@
Sinn Féin (SF) and the Democratic Unionist Party (DUP) nominated individuals for the First Minister and the deputy First Minister. The position of Justice Minister was awarded to the Alliance Party (AP) as part of a political compromise requiring cross-community support from both nationalists and unionists. The remaining positions were then allocated to the parties sequentially. For instance, the next position was allocated to SF because
@@
where the divisor for SF had been updated in a similar manner.
@@
$27/1$ >> $25/1$ >> $27/2$ >> $25/2$ >> $9/1$ $\geq$ $27/3$ >> $17/2$.
Although sequential apportionment is intended to streamline negotiations over ministerial positions, the Democratic Unionist Party postponed the formation of the Executive Committee, prompting legal intervention to avert a snap election. The process was not completed until 2024, suggesting that sequential apportionment did not, in this instance, speed the formation of the Executive Committee.
The use of the divisors $1,2,3,\ldots,$ distinguishes D’Hondt’s method. More generally, divisor methods can be defined through their unique set of denominators or divisors. In this article, we focus on the class of stationary divisor methods, which are parametrized by a fixed constant $c\in[0,1]$, and given by the sequence $a_{i}+c$ where $a_{i}$ is the nonnegative integer number of positions already awarded to party $i$, that is, party $i$’s current apportionment. We refer to the value $c$ as the cutpoint of the corresponding divisor method. The process of sequential apportionment for a general cutpoint $c$ is defined below.
@@
For fixed $0\leq c\leq 1$, given a distribution of votes $\mathbf{p}=(p_{1},p_{2},\ldots,p_{n})$, the apportionment sequence for the stationary divisor method with cutpoint $c$ is defined inductively as follows. Start by awarding each party $a_{i}=0$ seats. Once parties $p_{i}$ have been awarded $a_{i}$ seats, award the next seat to the party that maximizes the ratio $\frac{p_{i}}{a_{i}+c}$. Continue until all $h$ seats have been allocated.
If $c=1$, then the sequential apportionment process corresponds to the D’Hondt method. If $c=1/2$, the process corresponds to the Sainte-Laguë or Webster method. If $c=0$, the apportionment method is known as Adams’ method. Because dividing by 0 is undefined, initial adjustments are made for Adams’ method: we assume that $\frac{p_{i}}{0}>\frac{p_{j}}{0}$ whenever $p_{i}>p_{j}$ and $\frac{p_{i}}{0}>\frac{p_{j}}{k}$ for any $k>0$. Thus, under Adams’ method, the first $n$ seats are awarded, one to each party, assuming there are sufficient seats. There are other divisor methods, such as the Hill-Huntington method and Dean’s method, based on the geometric and harmonic means, respectively, that also give a seat to every party before a party gets a second seat. (To see how Hill-Huntington’s method fits into the sequential framework, it awards the next seat to the party that maximizes the ratio $p_{i}/\sqrt{a_{i}(a_{i}+1)}$, where $\sqrt{a_{i}(a_{i}+1)}$ is the geometric mean of the current apportionment $a_{i}$ and the next apportionment $a_{i}+1$.)
@@
Suppose two parties receive votes $\mathbf{p}=(32,14)$ with a house size of $h=10$. Both Adams and D’Hondt apportion $7$ of the $10$ seats to party 1 and $3$ seats to party 2. However, the inequalities used to generate the apportionment sequences show that the orders in which the seats are apportioned are different. In particular, the inequalities
$\text{Adams }c=0\qquad$ $\frac{32}{0}>\frac{14}{0}>\overbrace{\frac{32}{1}>\frac{32}{2}>\frac{14}{1}>\frac{32}{3}>\frac{32}{4}>\frac{14}{2}>\frac{32}{5}>\frac{32}{6}}^{\text{same sequence of inequalities}}$
$\text{D'Hondt }c=1\qquad$ $\underbrace{\frac{32}{1}>\frac{32}{2}>\frac{14}{1}>\frac{32}{3}>\frac{32}{4}>\frac{14}{2}>\frac{32}{5}>\frac{32}{6}}_{\text{same sequence of inequalities}}>\frac{14}{3}>\frac{32}{7}$
yield the apportionment sequences
[…]
Notice that the behavior after the first $2$ seats under Adams’ method matches the first $8$ seats under the D’Hondt method. This is generalized below. The sequence also encodes the apportionment for every house size less than 10; for example, if only $8$ seats were apportioned, the sequences indicate that party 1 would receive $5$ seats under Adams and $6$ seats under D’Hondt.
We focus on the apportionment sequences that include apportionments for all house sizes $h$ by considering the infinite sequence.
@@
For fixed $\mathbf{p}$, cutpoint $c$, and a house size $h$, let $S_{c}(h,\mathbf{p})=(s_{1},s_{2},\dots,s_{h})$ be the apportionment sequence under the stationary divisor method with cutpoint $c$ where $s_{j}=i$ if party $i$ receives the $j$th seat for each $j=1,\ldots,h$. Let $S_{c}(\mathbf{p})=\lim_{h\rightarrow\infty}S_{c}(h,\mathbf{p})$.
[…]
@@
For every positive vote vector $\mathbf{p}=(p_{1},p_{2},\ldots,p_{n})$, if $S_{1}(\mathbf{p})=(s_{1},s_{2},\dots)$ is the sequential apportionment under D’Hondt’s method, then the sequential apportionment under Adams’ method is $S_{0}(\mathbf{p})=(t_{1},t_{2},\dots)=(1,2,\dots,n,s_{1},s_{2},s_{3},\dots)$.
@@
We proceed by induction. First we show our base case that $t_{n+1}=s_{1}$. Under Adams’ method, each of the parties $1,2,\dots,n$ receive, in this order, one of the first $n$ seats, so that $t_{i}=i$ for $i=1$ to $n$. Seat $n+1$ is given to the party maximizing $\frac{p_{i}}{1}$, which is party 1 because $p_{1}\geq p_{i}$ for $i=2$ to $n$; hence, $t_{n+1}=1$. Under D’Hondt’s method, $s_{1}=1$ because $i=1$ maximizes the same ratio: $\frac{p_{i}}{1}$. And, $t_{n+1}=s_{1}$.
Assume that $t_{i+n}=s_{i}$ for $i=2$ to $k$, then we show that $t_{k+n+1}=s_{k+1}$. Let $\mathbf{a}$ be the apportionment under the D’Hondt method for $h=k$. Then, $a_{j}$ is the number of $s_{i}=j$ for $i=1$ to $k$. By definition, seat $k+1$ goes to the party with the smallest index $i$ among those indices that maximize $\frac{p_{i}}{a_{i}+1}$, or equivalently,
$\min_{i}\operatorname*{argmax}_{i=1\text{ to }n}\left\{\frac{p_{i}}{a_{i}+1}\right\}.$
By the inductive hypothesis, $b_{j}=a_{j}+1$ is party $j$’s apportionment under Adams’ method for $h=k+n$. It follows that seat $k+n+1$ goes to party
$\min_{i}\operatorname*{argmax}_{i=1\text{ to }n}\left\{\frac{p_{i}}{b_{i}}\right\}=\min_{i}\operatorname*{argmax}_{i=1\text{ to }n}\left\{\frac{p_{i}}{a_{i}+1}\right\}.$
Hence, $t_{k+n+1}=s_{k+1}$, completing the proof. ∎
Many properties of apportionment methods are best described in terms of the number of seats each party receives. Let $F_{c}(h,\mathbf{p})=\mathbf{a}=(a_{1},a_{2},\ldots,a_{n})$ indicate that after $h$ seats, the apportionment method has allocated $a_{i}$ seats to party $p_{i}$, where $h=\sum_{i}a_{i}$. Thus, in Example 2, $F_{0}(10,\mathbf{p})=F_{1}(10,\mathbf{p})=(7,3)$. The relationship between the apportionments from the Adams and the D’Hondt methods from Theorem 1 can be recast in terms of the number of seats apportioned.
Corollary 1.
For every positive vote vector $\mathbf{p}=(p_{1},p_{2},\ldots,p_{n})$, if $F_{1}(h,\mathbf{p})=\mathbf{a}$, then $F_{0}(h+n,\mathbf{p})=\mathbf{a}+\mathbf{1}$ where $\mathbf{1}$ is the vector of all 1s.
@@ 2.1 A Refinement of the Notion of Bias
The tendency of an apportionment method to systematically favor larger or smaller parties relative to exact proportionality is referred to as size bias. Although bias was part and parcel of the debate over the use of methods to apportion the U.S. House of Representatives in the 19th century, Huntington [10, 11] compared different methods in terms of how they optimize different measures of error from exact proportionality. Balinski and Young [2] compared divisor methods in terms of bias, providing the following definition and proposition.
@@
If $F_{c}$ and $F_{c^{\prime}}$ are stationary divisor methods corresponding to cutpoints $c$ and $c^{\prime}$, respectively, with $c>c^{\prime}$, then $F_{c}$ favors large parties relative to $F_{c^{\prime}}$.
The proposition implies that, among stationary methods, the D’Hondt method is the most biased in favor of large parties, while Adams’ method is the most biased in favor of small parties. Indeed, it is well-known that these two methods are the most biased in terms of size among all divisor methods, not just among the class of stationary methods.
@@
Suppose $c>c^{\prime}$ and that there exists $\mathbf{p}$ and an $h$ such that $S_{c}(h,\mathbf{p})>S_{c^{\prime}}(h,\mathbf{p})$. Let $h^{\prime}+1$ be the first house size for which $S_{c}(h^{\prime}+1,\mathbf{p})$ and $S_{c^{\prime}}(h^{\prime}+1,\mathbf{p})$ differ. Thus
@@
$cp_{i}+c^{\prime}p_{j}>cp_{j}+c^{\prime}p_{i}\iff(c-c^{\prime})(p_{i}-p_{j})>0.$
Since $c>c^{\prime}$, this implies $p_{i}>p_{j}$. But, due to the ordering of the parties based on the number of votes, then $i<j$. This contradiction proves the claim. ∎
By Proposition 2, for every $h$, $S_{0}(h,\mathbf{p})\geq S_{c}(h,\mathbf{p})\geq S_{1}(h,\mathbf{p})$ for all $c\in(0,1)$. Since this holds for all $h$, it follows that $S_{0}(\mathbf{p})\geq S_{c}(\mathbf{p})\geq S_{1}(\mathbf{p})$. This lexicographical ordering of apportionment sequences offers a refined perspective on bias, showing how bias is less an absolute statement of the number of seats a party receives and more a reflection of when large or small parties receive their seats.
To achieve this refinement of bias, we need to recall additional properties (homogeneity and weak proportionality) of divisor methods. Divisor methods satisfy homogeneity because the apportionment only depends on the relative proportion of votes $p_{i}/P$ where $P=p_{1}+\cdots+p_{n}$, noticeable in the definition of sequential apportionment.
Definition 4.
An apportionment method $F$ satisfies homogeneity if and only if $F(h,\mathbf{p})=F(h,\lambda\mathbf{p})$ for any $\lambda>0$.
Definition 5.
Let $\mathbf{a}=(a_{1},\ldots,a_{n})\in\mathbb{N}^{n}$, $p_{i}=\lambda a_{i}$ for all $i$, and $h=\sum_{i}a_{i}$. An apportionment method $F$ satisfies weak proportionality if and only if $F(h,\mathbf{p})=\mathbf{a}$.
In words, weak proportionality states that if $\mathbf{a}$ is proportional to the vote vector $\mathbf{p}$, then $\mathbf{a}$ should be the outcome of the apportionment method for $h=\sum_{i}a_{i}$. All divisor methods satisfy weak proportionality, which is illustrated with a continuation of Example 2.
@@
Let $\mathbf{p}=(32,14)$. Since $\mathbf{a}=(16,7)$ and $\mathbf{a}=(32,14)$ are both proportional to $\mathbf{p}$, then under D’Hondt’s method, $F_{1}(23,\mathbf{p})=(16,7)$ and $F_{1}(46,\mathbf{p})=(32,14).$ Additionally, their corresponding sequences are
@@
The proof that divisor methods satisfy weak proportionality follows most easily from an equivalent (non-sequential) definition of divisor methods. Under this definition, each apportionment method corresponds to a “rounding rule” $f$ that assigns to each positive integer $k$ a value $f(k)\in[k-1,k]$. Numbers in [k1,f(k))[k-1,f(k)) are rounded down to k1k-1 and numbers in $[f(k),k]$ are rounded up to $k$. (The function must also satisfy $f(k)\neq f(k+1)$, which could only happen if both were $k$.) The divisor method associated with rounding rule $f$ is defined as follows.
@@
A consequence of weak proportionality is that there are an infinite number of instances of $h$ (such as multiples of $\mathbf{p}$) where the apportionment agrees for all cutpoints $c$. All stationary methods are in synch at these values of $h$. Our perspective of bias is that it gives insight as to what happens between these values of commonality.